# Parameter uncertainties for imperfect surrogate models in the low-noise regime

### v3

@@ Abstract
Bayesian regression determines model parameters by minimizing the expected loss, an upper bound to the true generalization error. However, the loss ignores misspecification, where models are imperfect. Parameter uncertainties from Bayesian regression are thus significantly underestimated and vanish in the large data limit. This is particularly problematic when building models of low-noise, or near-deterministic, calculations, as the main source of uncertainty is neglected. We analyze the generalization error of misspecified, near-deterministic surrogate models, a regime of broad relevance in science and engineering. We show posterior distributions must cover every training point to avoid a divergent generalization error and design an ansatz that respects this constraint, which for linear models incurs minimal overhead. This is demonstrated on model problems before application to thousand dimensional datasets in atomistic machine learning. Our efficient misspecification-aware scheme gives accurate prediction and bounding of test errors where existing schemes fail, allowing this important source of uncertainty to be incorporated in computational workflows.
Surrogate models are widely used across science and engineering to efficiently approximate the action of computationally intensive simulation engines  Alizadeh et al. (2020); Deringer et al. (2019); Lapointe et al. (2020); Nyshadham et al. (2019); Montes de Oca Zapiain et al. (2022); Bonatti et al. (2022); Kudela and Matousek (2022). Domain expertise is leveraged to design optimal features and model architecture before parameters are inferred from training data generated by the engines. In a broad range of applications, the inference problem shares three key characteristics:
a) Simulation engines are near deterministic: outputs have vanishing aleatoric uncertainty, e.g. the atomic energy $V({\bf X})$ is deterministic to atomic positions ${\bf X}$.
b) Surrogate models are misspecified: no single choice of the $P$ parameters can match all $N$ observations, meaning parameters are intrinsically uncertain.
c) Surrogate models are underparametrized: large quantities of training data are available, such that standard epistemic uncertainty estimates from Bayesian inference vanish, typically corresponding to the regime $P/N\ll 1$.
@@
In this paper, we analyze the generalization error of deterministic, misspecified and under-parametrized surrogate models, i.e. those typically used to approximate simulation engines in science and engineering. As posterior distributions strictly become unbounded in the deterministic limit, this is only taken as a theoretical device to derive conditions that any posterior must obey to avoid a divergent generalization error. We then design an ansatz that respect this condition and admits a robust learning scheme at finite $N$, with standard treatment of epistemic ($N$-dependent) errors, reducing to a constrained loss minimization problem for large $N/P$. Our final misspecification-aware parameter posterior has finite uncertainty even as $N/P\to\infty$ with a provably lower generalization error than the Gibbs posterior of Bayesian inference.
@@
a) We define pointwise optimal parameter sets (POPS) for each training point, within which model predictions are exact. Parameter distributions must have mass in every POPS to avoid a divergent generalization error. b) We use the ensemble of loss minimizers from each POPS to design two ansatz posteriors, a reweighted $\mathcal{O}(N)$ ensemble and a bounding $\mathcal{O}(P)$ hypercube, both of which respect POPS occupancy and give a finite generalization error, lower than the Gibbs posterior from inference.
@@
c) For linear models, our ansatz can be efficiently evaluated via rank-one updates to a leverage-weighted loss minimizer. This gives an efficient scheme to incorporate misspecification into linear regression with minimal computational overhead.
@@
d) We test the approach on synthetic datasets, showing how our scheme provides robust bounds on test errors. This performance is maintained under application to thousand-dimensional datasets from atomic machine learning, untreatable with existing methods.
.1 Related work
Germain and coworkers Germain et al. (2016) established an important connection between Bayesian inference and probability approximately correct (PAC-Bayes) analysis for regression problems with unbounded losses, which bounds true expectations by empirical likelihood expectations through concentration inequalities. This provides a robust learning scheme without test-train splitting, and gives a rationalization for Bayesian inference as a form of regularization. In particular, minimizing the PAC-Bayes bound for the expected loss gives the familiar relation between posterior, prior and likelihood from Bayesian inference.
In recent years, multiple groups have used the PAC-Bayes framework to derive tighter bounds for the generalization error than that provided by the expected loss  Masegosa (2020); Germain et al. (2016); Lahlou et al. (2021); Psaros et al. (2023); Lotfi et al. (2022); Morningstar et al. (2022). However, these efforts focus on probabilistic regression settings, often with neural network models far from the underparametrized limit, meaning aleatoric, epistemic and misspecification errors must be considered jointly.
Masegosa Masegosa (2020) derived second order PAC-Bayes bounds for the generalization error minimized through a variational or ensemble ansatz. Morningstar et al. Morningstar et al. (2022) developed estimators to assess disagreement between the generalization error and expected loss, deriving specialized PAC bounds for theoretical guarantees. Lahlou et al. Lahlou et al. (2021) considered misspecification of neural networks, jointly training a minimum loss surrogate and an independent predictor of misspecification error. Lofti et al. Lotfi et al. (2022) considered misspecification in the context of deep model selection, deriving conditional likelihoods better aligned with the generalization error.
As various definitions for the distinction between epistemic and misspecification uncertainties exist Kato et al. (2022), we define misspecification as non-aleatoric uncertainties which survive in the underparametrized limit, where the minimizer of PAC-Bayes bounds for the expected loss, which coincides with the posterior from Bayesian inference, predict vanishing parameter uncertainties. This definition is made more precise in section I.4. Our epistemic uncertainties thus coincide with standard estimates from loss minimization.
I Bayesian surrogate models
@@ I.1 Deterministic simulation engines
A simulation engine ${\mathcal{E}}:{\mathcal{X}}\to{\mathcal{Y}}$ takes input ${\mathbf{X}}\in{\mathcal{X}}$ and returns ${\mathbf{Y}}\in{\mathcal{Y}}$. We consider the regime where the aleatoric uncertainty of ${\mathcal{E}}$ is weak and can be described as a Gaussian distribution with mean $\langle{\mathbf{Y}}|{\mathbf{X}}\rangle\equiv{\mathcal{E}}({\mathbf{X}})$ and covariance $\langle{\mathbf{Y}}{\mathbf{Y}}^{\top}|{\mathbf{X}}\rangle-{\mathcal{E}}({\mathbf{X}}){\mathcal{E}}^{\top}({\mathbf{X}})\equiv{\bm{\Sigma}}_{\mathcal{Y}}({\mathbf{X}})$, where $\|{\bm{\Sigma}}_{\mathcal{Y}}({\mathbf{X}})\|=\mathcal{O}(\epsilon^{2Y})$, and $\epsilon$ is assumed small. The true conditional output distribution thus reads
$\rho_{\mathcal{E}}({\mathbf{Y}}|{\mathbf{X}})=\mathcal{N}({\mathbf{Y}}|{\mathcal{E}}({\mathbf{X}}),{\bm{\Sigma}}_{\mathcal{Y}}({\mathbf{X}})).$ (1)
We assume homoskedatiscity ${\bm{\Sigma}}_{\mathcal{Y}}({\mathbf{X}})={\bm{\Sigma}}_{\mathcal{Y}}$ of this weak aleatoric error to investigate the deterministic, underparametrised limit $\epsilon\to 0$, $N/P\to\infty$ for $\epsilon=\mathcal{O}(\sqrt{P/N})$. With weights $\sum_{i=1}^{N}{\rm w}_{i}=1$, the training configurations ${\mathcal{D}}_{N}=\{{\mathbf{Y}}_{i},{\mathbf{X}}_{i}\}_{i=1}^{N}$ form an input distribution in ${\mathbf{X}}$ of
@@
Learning requires that we can bound expectations over independent identically distributed (i.i.d.) samples ${\mathbf{X}},{\mathbf{Y}}\in{\mathcal{D}}$
@@
For timeseries or other signal data the i.i.d. property is only satisfied for decorrelated segments ${\mathbf{X}}=\{{\bf x}_{t}\}_{t=1}^{T}$, meaning ${\bm{\Sigma}}_{\mathcal{Y}}({\mathbf{X}})$ captures signal correlations as in generalized least squares regression.
@@ I.2 Deterministic surrogate models
Unlike typical Bayesian schemes, we do not treat the covariance term ${\bm{\Sigma}}_{\mathcal{Y}}$ as fitting parameter but as an intrinsic property of ${\mathcal{E}}$. This has implications for model selection criteria, discussed in the supplementary material (SM) Swinburne and Perez (2024).
@@ I.3 The generalization error and expected loss
In principle any learning scheme should aim to minimize the non-convex functional ${\mathcal{G}}[\pi]$, but in practice learning schemes target a convex upper bound, the expected loss ${\mathcal{L}}[\pi]\geq{\mathcal{G}}[\pi]$, often known as the negative log likelihood. Using the Jensen inequality $-\langle\ln x\rangle\leq-\ln\langle x\rangle$ gives
@@
${\mathcal{L}}[\pi]$ is clearly minimized by a delta function $\delta({\bm{\Theta}}-{\bm{\Theta}}^{*}_{\mathcal{L}})$, where ${\bm{\Theta}}^{*}_{\mathcal{L}}\in\arg\min_{\mathcal{P}}\left\langle-\ln\left|\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})\right|\right\rangle_{\mathcal{D}}$, i.e. vanishing parameter uncertainty. The next section shows how this emerges from the inference posterior at finite $N/P$.
@@ I.4 PAC-Bayes estimation and connection to inference
PAC-Bayes analysis Alquier (2021) provides concentration inequalities Hoeffding (1994) that bound expectations over ${\mathcal{D}}$ by empirical expectations over ${\mathcal{D}}_{N}$ with some probability. Whilst originally designed for classification problems, the approach has been extended to regression settings with unbounded losses Germain et al. (2016). As discussed in the SM Swinburne and Perez (2024), The expected loss ${\mathcal{L}}[\pi]$ admits the PAC-Bayes upper bound, holding with probability $1-\xi$, of
@@
where $\hat{L}_{N}({\bm{\Theta}})=-\langle\ln|\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})\rangle_{N}$ is the empirical loss over $\rho_{N}({\mathbf{X}})$, $\pi_{0}({\bm{\Theta}})$ is some prior and $C_{0}$ is a constant Swinburne and Perez (2024). The minimizer of the upper bound (21) is the posterior from Bayesian inference Germain et al. (2016)
@@
As $N/P\to\infty$ the posterior is strongly peaked around a loss minimizer ${\bm{\Theta}}^{*}_{\mathcal{L}}\in\arg\min_{{\bm{\Theta}}\in{\mathcal{P}}}L({\bm{\Theta}})$. Defining ${\bm{\Sigma}}^{-1}_{0}$ as the curvature of $-\ln\pi_{0}({\bm{\Theta}})$ around ${\bm{\Theta}}^{*}_{\mathcal{L}}$, we use Laplace’s method Swinburne and Perez (2024) as $N/P\to\infty$ to write

### v5

@@ Abstract
Bayesian regression determines model parameters by minimizing the expected loss, an upper bound to the true generalization error. However, this loss ignores model form error, or misspecification, meaning parameter uncertainties are significantly underestimated and vanish in the large data limit. As misspecification is the main source of uncertainty for surrogate models of low-noise calculations, such as those arising in atomistic simulation, predictive uncertainties are systematically underestimated. We analyze the true generalization error of misspecified, near-deterministic surrogate models, a regime of broad relevance in science and engineering. We show that posterior parameter distributions must cover every training point to avoid a divergence in the generalization error and design a compatible ansatz which incurs minimal overhead for linear models. The approach is demonstrated on model problems before application to thousand-dimensional datasets in atomistic machine learning. Our efficient misspecification-aware scheme gives accurate prediction and bounding of test errors in terms of parameter uncertainties, allowing this important source of uncertainty to be incorporated in multi-scale computational workflows.
Surrogate models are widely used across science and engineering to efficiently approximate the output of computationally intensive simulation engines  Alizadeh et al. (2020); Deringer et al. (2019); Lapointe et al. (2020); Nyshadham et al. (2019); Montes de Oca Zapiain et al. (2022); Bonatti et al. (2022); Kudela and Matousek (2022). Domain expertise is leveraged to design optimal features and model architecture before parameters are inferred from training data generated by the engines. In a broad range of applications, the inference problem shares three key characteristics:
  1. a)
    Simulation engines are near deterministic: outputs have vanishing aleatoric uncertainty, e.g. the atomic energy $V({\bf X})$ from some quantum chemistry calculation is a near-deterministic function of atomic positions ${\bf X}$, with very weak uncertainty due to e.g. convergence of the electronic minimization scheme.
  2. b)
    Surrogate models are misspecified: no single choice of the $P$ parameters can match all $N$ observations, meaning parameters are intrinsically uncertain.
  3. c)
    Surrogate models are underparametrized: large quantities of training data are available, such that standard epistemic uncertainty estimates from Bayesian inference vanish, typically corresponding to the regime $P/N\ll 1$.
@@
In this paper, we analyze the generalization error of near-deterministic, misspecified and under-parametrized surrogate models, i.e. those typically used to approximate simulation engines in science and engineering. Our central application is to the problem of fitting interatomic potentials for atomic simulationsGoryaeva et al. (2021); Montes de Oca Zapiain et al. (2022); Li et al. (2020); Wood and Thompson (2018), where the training data is generated by quantum chemistry calculations. However, our analysis applies to the fitting of any surrogate model to a near-deterministic ground truth. Posterior distributions strictly become unbounded as the aleatoric uncertainty vanishes, meaning the low-noise/near-deterministic limit is only taken as a theoretical device to derive conditions that any posterior distribution must obey to avoid a logarithmic divergence in the generalization error of that distribution. We then design an ansatz that respect this condition and admits a robust learning scheme at finite $N$, with standard treatment of epistemic ($N$-dependent) errors, reducing to a constrained loss minimization problem for large $N/P$. Our final misspecification-aware parameter posterior has finite uncertainty even as $N/P\to\infty$, whilst the generalization error of the posterior from Bayesian inference diverges. The approach thus captures misspecification uncertainties missed by standard Bayesian regression schemes.
Our implementation for linear models is a robust and to our knowledge unique means to estimate misspecification uncertainties, and thus is the optimal available choice when fitting high-dimensional linear models to low-noise data.
@@
    We define pointwise optimal parameter sets (POPS) for each training point, within which model predictions are exact. Parameter distributions must have mass in every POPS to avoid a divergent generalization error.
  2. b)
    We use the ensemble of loss minimizers from each POPS to design two ansatz posteriors, a reweighted $\mathcal{O}(N)$ ensemble and a bounding $\mathcal{O}(P)$ hypercube, both of which respect POPS occupancy and give a finite generalization error, lower than the posterior from Bayesian inference.
  3. c)
@@
    For linear models, our ansatz can be efficiently evaluated via rank-one updates to a leverage-weighted loss minimizer. This gives an efficient scheme to incorporate misspecification into linear regression, which we show gives minimal computational overhead (approximately a factor of two in training) over Bayesian inference.
  4. d)
    A pip-installable Python implementation of our misspecification-aware POPS regression algorithm, following the Scikit-learn linear_model API Pedregosa et al. (2011), is provided at https://github.com/tomswinburne/POPS-Regression.git.
  5. e)
@@
    We test the approach on synthetic datasets, showing how our scheme provides robust bounds on test errors. This performance is maintained under application to thousand-dimensional datasets for fitting machine learning interatomic potentialsMontes de Oca Zapiain et al. (2022); Li et al. (2020) and informatics modelsTynes et al. (2024); Ruddigkeit et al. (2012), untreatable with existing misspecification-aware methods.
The remainder of this paper is organized as follows. In Section I, we review Bayesian regression and PAC-Bayes bounds, highlighting the limitations of standard approaches for misspecified models. Section II presents our analysis of the generalization error for near-deterministic surrogate models, introducing the concept of pointwise optimal parameter sets (POPS). In Section III, we develop our misspecification-aware ansatz, providing efficient implementations for linear models. Section IV demonstrates the effectiveness of our approach on both synthetic datasets and high-dimensional problems in atomic machine learning. Section V discusses our method in the context of existing uncertainty quantification and propagation schemes for atomistic simulation. Finally, Section VI summarizes our findings and highlights future directions for research in misspecification-aware regression.
I Bayesian fitting of surrogate models
This section summarizes known results from Bayesian analysis and PAC-Bayes inequalitiesAlquier (2021), focusing on the limit of interest for this paper: fitting a misspecified surrogate model in the case that the ground truth is (near-)deterministic, i.e. known to have vanishing aleatoric error. We provide a glossary of all terms in Appendix A
@@ I.1 Deterministic simulation engines
A simulation engine ${\mathcal{E}}:{\mathcal{X}}\to{\mathcal{Y}}$ takes input ${\mathbf{X}}\in{\mathcal{X}}$ and returns ${\mathbf{Y}}\in{\mathcal{Y}}$. We denote the dimension of ${\mathcal{Y}}$ and ${\mathcal{X}}$ by $Y$ and $X$ respectively. We consider the regime where the aleatoric uncertainty of ${\mathcal{E}}$ is weak, such that the output ${\mathbf{Y}}$ follows a Gaussian distribution
$\rho_{\mathcal{E}}({\mathbf{Y}}|{\mathbf{X}})=\mathcal{N}({\mathbf{Y}}|{\mathcal{E}}({\mathbf{X}}),{\bm{\Sigma}}_{\mathcal{Y}}({\mathbf{X}})),$ (1)
which is fully defined by the output mean $\langle{\mathbf{Y}}|{\mathbf{X}}\rangle\equiv{\mathcal{E}}({\mathbf{X}})$ and covariance $\langle{\mathbf{Y}}{\mathbf{Y}}^{\top}|{\mathbf{X}}\rangle-{\mathcal{E}}({\mathbf{X}}){\mathcal{E}}^{\top}({\mathbf{X}})\equiv{\bm{\Sigma}}_{\mathcal{Y}}({\mathbf{X}})$, where $\|{\bm{\Sigma}}_{\mathcal{Y}}({\mathbf{X}})\|=\mathcal{O}(\epsilon^{2Y})$, and $\epsilon$ is assumed small. As the aleatoric error is vanishingly weak, we assume homoskedatiscity ${\bm{\Sigma}}_{\mathcal{Y}}({\mathbf{X}})={\bm{\Sigma}}_{\mathcal{Y}}$ for simplicity. We will later take the deterministic, underparametrised limit $\epsilon\to 0$, $N/P\to\infty$ assuming that $\epsilon=\mathcal{O}(\sqrt{P/N})$. With weights $\sum_{i=1}^{N}{\rm w}_{i}=1$, the training data ${\mathcal{D}}_{N}=\{{\mathbf{Y}}_{i},{\mathbf{X}}_{i}\}_{i=1}^{N}$ form an input distribution in ${\mathbf{X}}$ of
@@
Learning requires that we can bound expectations over independent identically distributed (i.i.d.) samples ${\mathbf{X}},{\mathbf{Y}}\in{\mathcal{D}}$, where ${\mathcal{D}}$ is the set of all possible input-output pairs, corresponding to the limit of infinite training data,
@@
For timeseries or other signal data the i.i.d. property is only satisfied for decorrelated segments ${\mathbf{X}}=\{{\bf x}_{t}\}_{t=1}^{T}$, meaning ${\bm{\Sigma}}_{\mathcal{Y}}$ should capture signal correlations as in generalized least squares regressionAlquier (2021); Germain et al. (2016); Masegosa (2020).
@@ I.2 Deterministic surrogate models
Unlike typical Bayesian schemesBhat and Kumar (2010); Von Toussaint (2011), we do not treat the covariance term ${\bm{\Sigma}}_{\mathcal{Y}}$ as fitting parameter but as an intrinsic property of ${\mathcal{E}}$. This has implications for model selection criteria, discussed in appendix B.
@@ I.3 The generalization error and expected loss
In principle any learning scheme should aim to minimize the functional ${\mathcal{G}}[\pi]$, but in practice learning schemes target a tractable upper bound, the expected loss ${\mathcal{L}}[\pi]\geq{\mathcal{G}}[\pi]$, often known as the negative log likelihood. Using the Jensen inequality $-\langle\ln x\rangle\leq-\ln\langle x\rangle$, as used to derive Gibbs’ famous inequality for free energy estimationGibbs (1902), we find that
@@
${\mathcal{L}}[\pi]$ is clearly minimized by a delta function $\delta({\bm{\Theta}}-{\bm{\Theta}}^{*}_{\mathcal{L}})$, where ${\bm{\Theta}}^{*}_{\mathcal{L}}\in\arg\min_{\mathcal{P}}\left\langle-\ln\left|\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})\right|\right\rangle_{\mathcal{D}}$. As the delta function has formally zero width, the loss minimizing distribution thus has vanishing parameter uncertainty. The next section shows how this emerges from the inference posterior at finite $N/P$.
@@ I.4 PAC-Bayes estimation and connection to inference
PAC-Bayes analysis Alquier (2021) provides concentration inequalities Hoeffding (1994) that bound expectations over ${\mathcal{D}}$ by empirical expectations over ${\mathcal{D}}_{N}$ with some probability. Whilst originally designed for classification problems, Germain and coworkers Germain et al. (2016) established an important connection between Bayesian inference and probability approximately correct (PAC-Bayes) analysis for regression problems with unbounded losses, which bounds true expectations by empirical likelihood expectations through concentration inequalities. A short additional note on the PAC-Bayes framework for regression is provided in appendix LABEL:app:pac_constant.
The PAC-Bayes framework provides a robust learning scheme without test-train splitting, and gives a rationalization for Bayesian inference as a form of regularization. In particular, minimizing the PAC-Bayes bound for the expected loss gives the familiar relation between posterior, prior and likelihood from Bayesian inference. This subsection adapts the results of Germain et al. Germain et al. (2016) and Masegosa et al. Masegosa (2020) in the near-deterministic limit of interest here. The central result ofGermain et al. (2016) is that the expected loss ${\mathcal{L}}[\pi]$ admits the PAC-Bayes upper bound, holding with probability $1-\xi$, of
@@
where $\hat{L}_{N}({\bm{\Theta}})=-\langle\ln|\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})\rangle_{N}$ is the empirical loss over $\rho_{N}({\mathbf{X}})$, $\pi_{0}({\bm{\Theta}})$ is some prior and $C_{0}$ is a constant. The value of the constant $C_{0}$ requires certain assumptions on the distribution of the error between the empirical and true expected loss. Following (Germain et al., 2016), we assume that this error distribution is unbounded but sub-normal, i.e. all error fluctuations are bounded by a normal distribution of variance $s^{2}$, giving $C_{0}[\pi_{0}]<s^{2}/2$. We refer the reader to (Germain et al., 2016) and subsequent works (Shalaeva et al., 2020; Masegosa, 2020) for further discussion. The minimizer of the upper bound (8) is the posterior from Bayesian inference Germain et al. (2016), reading, with a parameter $\lambda$ to enforce normalization
@@
As $N/P\to\infty$ the posterior is strongly peaked around a loss minimizer ${\bm{\Theta}}^{*}_{\mathcal{L}}\in\arg\min_{{\bm{\Theta}}\in{\mathcal{P}}}L({\bm{\Theta}})$. Defining ${\bm{\Sigma}}^{-1}_{0}$ as the curvature of $-\ln\pi_{0}({\bm{\Theta}})$ around ${\bm{\Theta}}^{*}_{\mathcal{L}}$, we use Laplace’s method (appendix C) as $N/P\to\infty$ to write
@@
I.5 Related work on misspecification-aware Bayesian regression
In recent years, multiple groups have used the PAC-Bayes framework to derive tighter bounds for the generalization error than that provided by the expected loss  Masegosa (2020); Germain et al. (2016); Lahlou et al. (2021); Psaros et al. (2023); Lotfi et al. (2022); Morningstar et al. (2022), to build a misspecified-aware regression scheme. However, these efforts focus on probabilistic regression settings, often with neural network models far from the underparametrized limit, meaning aleatoric, epistemic and misspecification errors must be considered jointly.
Masegosa Masegosa (2020) derived second order PAC-Bayes bounds for the generalization error minimized through a variational or ensemble ansatz. Morningstar et al. Morningstar et al. (2022) developed estimators to assess disagreement between the generalization error and expected loss, deriving specialized PAC bounds for theoretical guarantees. Lahlou et al. Lahlou et al. (2021) considered misspecification of neural networks, jointly training a minimum loss surrogate and an independent predictor of misspecification error. Lofti et al. Lotfi et al. (2022) considered misspecification in the context of deep model selection, deriving conditional likelihoods better aligned with the generalization error. As various definitions for the distinction between epistemic and misspecification uncertainties exist Kato et al. (2022), we have defined (section I.4) misspecification as non-aleatoric uncertainties which survive in the underparametrized limit, where the minimizer of PAC-Bayes bounds for the expected loss, which coincides with the posterior from Bayesian inference, predicts vanishing parameter uncertainties. Our epistemic uncertainties thus coincide with standard estimates from loss minimization.